Calculate the wavelength (in nm) of the light radiation capable of causing an ionization energy equal to 502*J.𝑚𝑜𝑙−1

(c= 3*m/s;     h= 6.626 × J.)


1
Expert's answer
2021-04-07T11:47:55-0400

"E= hv"


"502 = (6.626\u00d710^{-34})v"


"v= \frac{502}{6.626×10^{-34}}=7.577×10^{35}"


"\lambda v= c"


"\lambda= \frac{c}{v}"


"\lambda=\frac{3.00×10^8}{7.577×10^{35}}=3.959×10^{42}"